East Dome Improvements
Historical changes, milestones and upgrades in the East Dome:
- 1989 - Club raises $23,000 to insure East Dome built at new George Observatory.
- May 1990 - HMNS loans FBAC 8" telescope for East Dome.
- March 1991 - Club buys 10" Meade scope as a temporary until the 16" is delivered.
- November 1991 - 10" Meade installation complete.
- August 1993 - FBAC purchases 18" Newtonian from Loyd Overcash for $5000.
- October 1993 - 18" Newtonian installed in East Dome.
- December 1993 - Member Steve Dessens purchases old 10" Meade from ED.
- March 1994 - New 80mm finder purchased for 18" Newtonian.
- May 1994 - New raised floor completed in East Dome by Mike Morton.
- September 1994 - Used PC purchased from Don Edwards for East Dome.
- May 1995 - New heavier RA shaft and bearings for 18" Newtonian.
- February 1996 - Club CookBook CCD camera completed by Keith Rivich.
- February 1997 - New 20" RA gear ordered from OptiCraft.
- March 1997 - Quadrant drive system and stepper motors ordered.
- January 1998 - Problems noted with OptiCraft and Quadrant installations.
- March 1999 - New Byers gear received to replace OptiCraft gear.
- August 2001 - New Quadrant GoTo drive system operational.
- February 2002 - East Dome Committee formed by President Keith Rivich.
- May 2002 - First beta copy of new East Dome operators manual by Jim Ellis.
- August 2002 - New computer donated by Calvin Embry installed in East Dome.
- August 2002 - Mirror cleaning.
- November 2002 - SBIG ST7 CCD camera purchased to replace the old Cookbook camera.
- August 2004 - Quadrant 4 purchased through donations by ED users but it has problems.
- August 2004 - UPS for East Dome shutter donated by Joe Dellinger.
- September 2004 - Dome rotation rollers replaced.
- May 2005 - First Aid kit purchased and installed in East Dome.
- May 2005 - Cynthia Gustava donated 17" computer monitor for East Dome.
- July 2005 - Shutter rollers replaced on East Dome.
- November 2005 - Foam skirt replaced around azimuth dome rollers.
- April 2006 - RA pillow block and CamRol bearings replaced on fork mount.
- April 2006 - Shutter rollers replaced again on East Dome.
- February 2007 - Quadrant 4 back from AIM Controls and working. Failed again mid-year.
- September 2007 - Club approves purchase of replacement drive for 18" in East Dome.
- November 2007 - Computer donated by Mark Holdsworth of Netslyder and installed.
- November 2007 - New scope cover built and installed.
- November 2007 - Sidereal Technology servo drive and Argo Navis DSC ordered.
- February 2008 - Sidereal Technology servo drive system installed and tested.
- March 2008 - New data conduits installed and cabling run from back of telescope pier to PCs.
- April 2008 - 15" LCD monitor for west side PC donated by Steve Clayworth.
- June 2011 - East Dome remodeling completed - See summary and photos.
Planned changes, upgrades, and maintenance in the East Dome:
- Replacement of wall paneling in dome.